OpenStudy (anonymous):

what are the zeros for x^4-6x^3+16x^2-24x+16

OpenStudy (anonymous):

\[x^4-6 x^3+16 x^2-24 x+16=(x-2)^2 \left(x^2-2 x+4\right) \]x = 2, 2 The other two zeros can be obtained by solving\[x^2-2 x+4=0 \]for x.\[\left\{x=1-i \sqrt{3},x=1+i \sqrt{3}\right\} \]